# Regularization by regular noise: a numerical result

### v1

@@ 4. Strong convergence rate
$\leqslant N(r-s_{1})^{1+\alpha H}\left\|\|B_{r}^{H}-B^{H}_{k_{n}(r)}\|_{L_{\omega}^{2}|\mathcal{F}_{s_{1}}}\right\|_{L^{p}_{\omega}}$
$=N(r-s_{1})^{1+\alpha H}\|{\mathbb{E}}^{s_{1}}|B_{r}^{H}-B_{k_{n}(r)}^{H}|^{2}\|^{\frac{1}{2}}_{L_{\omega}^{\frac{p}{2}}}$
$\leqslant N(r-s_{1})^{1+\alpha H}\|B_{r}^{H}-B^{H}_{k_{n}(r)}\|_{L^{p}_{\omega}}$
$\leqslant\frac{N}{n}(r-s_{1})^{1+\alpha H}.$
@@
$\|V_{1}\|_{L^{p}_{\omega}}$ $\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{5}}(r-s_{3})^{-(2-\alpha)H}(r-s_{1})^{1+\alpha H}\mathrm{d}r\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{2+(2\alpha-2)H}.$
@@
Since Assumption˜2.1 implies $2+(2\alpha-2)H>1$, we can obtain that (3.2) holds with $C_{2}=\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}$.
In order to show that the process $\mathcal{A}$ in (3.3) is given by
$\mathcal{A}_{t}=\int_{0}^{t}b(B^{H}_{r}+\varphi^{n}_{k_{n}(r)})-b(B^{H}_{k_{n}(r)}+\varphi_{k_{n}(r)}^{n})\mathrm{d}r,$
it suffices to show (3.4). For doing so, we first write
$\mathcal{A}_{t}-\mathcal{A}_{s}-A_{s,t}$ $=\int_{s}^{t}b(B_{r}^{H}+\varphi^{n}_{k_{n}(r)})-\mathcal{P}_{r-s_{1}}^{H}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)})\mathrm{d}r$
$\quad-\int_{s}^{t}b(B^{H}_{k_{n}(r)}+\varphi^{n}_{k_{n}(r)})-\mathcal{P}^{H}_{(k_{n}(r)-s_{1})\vee 0}b({\mathbb{E}}^{s_{1}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)})\mathrm{d}r.$
Then similar to what we have done in Lemma˜4.2 (that is, (4), (4.16) and (4.17)), we can conclude the claim by the uniqueness from Lemma˜3.1.
@@
$\|(X-X^{n})_{t}-(X-X^{n})_{s}\|_{L^{p}_{\omega}}$ $\leqslant N(\|\varphi-\varphi^{n}\|_{C^{\frac{1}{2}}_{p}}+n^{-1})|t-s|^{\frac{1}{2}+\varepsilon}$
$=N(\|X-X^{n}\|_{C^{\frac{1}{2}}_{p}}+n^{-1})|t-s|^{\frac{1}{2}+\varepsilon},$
@@
$[X-X^{n}]_{C_{p}^{\frac{1}{2}}[S,T]}\leqslant N(\|X-X^{n}\|_{C_{p}^{\frac{1}{2}}[S,T]}+n^{-1})(T-S)^{\varepsilon}.$
@@
$\|X-X^{n}\|_{C_{p}^{\frac{1}{2}}[S,T]}$ $\leqslant|(X-X^{n})_{S}|+2[X-X^{n}]_{C_{p}^{\frac{1}{2}}[S,T]}$
$\leqslant|(X-X^{n})_{S}|+N(\|X-X^{n}\|_{C_{p}^{\frac{1}{2}}[S,T]}+n^{-1})(T-S)^{\varepsilon}.$
@@ 5. Optimality
In the following we present a result concerning the optimal convergence rate of EM (1.9) for rather smooth $b\in{\mathcal{C}}^{1}$.
Theorem 5.1.
Suppose $b\in{\mathcal{C}}^{1}$ and $x_{0}^{n}=x_{0}$. Then for any $t\in[0,1]$, a.s.
$\lim_{n\rightarrow\infty}n(X_{t}-X_{t}^{n})=:c(t)$ (5.1)
exists and $c(t)$ satisfies
$c^{\prime}(t)=\nabla b(X_{t})c(t)+\frac{1}{2}\nabla b(X_{t})(b(X_{t})+(B_{t}^{H})^{\prime}),\quad c(0)=0.$
@@
$n(X_{t}-X_{t}^{n})$ $=\int_{0}^{t}n(b\left(X_{r}\right)-b\left(X_{r}^{n}\right))+n(b\left(X_{r}^{n}\right)-b(X_{k_{n}(r)}^{n}))\mathrm{d}r$
$=\int_{0}^{t}\int_{0}^{1}\nabla b(X_{r}^{n}+\theta(X_{r}-X_{r}^{n}))n(X_{r}-X_{r}^{n})$
$\qquad\quad+\nabla b(X_{k_{n}(r)}^{n}+\theta(X_{r}^{n}-X_{k_{n}(r)}^{n}))n(X_{r}^{n}-X_{k_{n}(r)}^{n})\mathrm{d}\theta\mathrm{d}r.$
By fundamental theory of calculus we can write
$n(X_{t}-X_{t}^{n})$ $=\int_{0}^{t}\exp\left(\int_{s}^{t}\int_{0}^{1}\nabla b(X_{r}^{n}+\theta(X_{r}-X_{r}^{n}))d\theta dr\right)$
$\qquad\int_{0}^{1}\nabla b(X_{k_{n}(s)}^{n}+\theta(X_{s}^{n}-X_{k_{n}(s)}^{n}))n(X_{s}^{n}-X_{k_{n}(s)}^{n})\mathrm{d}\theta\mathrm{d}s=:c^{n}(t).$
@@
Following from (1.9) and the fact a.s. $B^{H}\in{\mathcal{C}}^{H-\varepsilon}([0,1])$ for sufficiently small $\varepsilon>0$ by Kolmogorov continuity criteria, which means there exists a small enough $\epsilon^{\prime}>0$ so that a.s. $B_{H}\in{\mathcal{C}}^{1+\epsilon^{\prime}}$, we have a.s. (denote $\{c\}:=c-\lfloor c\rfloor$ for $c\in{\mathbb{R}}_{+}$)
$n(X_{s}^{n}-X_{k_{n}(s)}^{n})$ $=n(s-k_{n}(s))b(X_{k_{n}(s)}^{n})+n(B_{s}^{H}-B_{k_{n}(s)}^{H})$
$=\{ns\}b(X_{k_{n}(s)}^{n})+\{ns\}(B_{s}^{H})^{\prime}+o(n^{-\varepsilon^{\prime}});$

### v2

@@ 4. Strong convergence rate
$\leqslant C(r-s_{1})^{1+\alpha H}\left\|\|B_{r}^{H}-B^{H}_{k_{n}(r)}\|_{L_{\omega}^{2}|\mathcal{F}_{s_{1}}}\right\|_{L^{p}_{\omega}}$
$=C(r-s_{1})^{1+\alpha H}\|{\mathbb{E}}^{s_{1}}|B_{r}^{H}-B_{k_{n}(r)}^{H}|^{2}\|^{\frac{1}{2}}_{L_{\omega}^{\frac{p}{2}}}$
$\leqslant C(r-s_{1})^{1+\alpha H}\|B_{r}^{H}-B^{H}_{k_{n}(r)}\|_{L^{p}_{\omega}}$
$\leqslant\frac{C}{n}(r-s_{1})^{1+\alpha H}.$
@@
$\|V_{1}\|_{L^{p}_{\omega}}$ $\leqslant\frac{C}{n}\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{5}}(r-s_{3})^{-(2-\alpha)H}(r-s_{1})^{1+\alpha H}\mathrm{d}r\leqslant\frac{C}{n}\|f\|_{{\mathcal{C}}^{\alpha}}|t-s|^{2+(2\alpha-2)H}.$
@@
Since Assumption˜2.1 implies $2+(2\alpha-2)H>1$, we can obtain that (3.2) holds with $C_{2}=\frac{C}{n}\|b\|_{{\mathcal{C}}^{\alpha}}$.
Similarly to Lemma˜4.2, we could verify the process $\mathcal{A}$ in (3.3) is given by
$\mathcal{A}_{t}=\int_{0}^{t}f(B^{H}_{r}+\varphi^{n}_{k_{n}(r)})-f(B^{H}_{k_{n}(r)}+\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$
@@
$\|(X-X^{n})_{t}-(X-X^{n})_{s}\|_{L^{p}_{\omega}}$ $\leqslant C(\|\varphi-\varphi^{n}\|_{C^{\gamma}_{p}[s,t]}+n^{-1})|t-s|^{\gamma+\varepsilon}$
$=C(\|X-X^{n}\|_{C^{\gamma}_{p}[s,t]}+n^{-1})|t-s|^{\gamma+\varepsilon},$
@@
$[X-X^{n}]_{C_{p}^{\gamma}[S,T]}\leqslant C(\|X-X^{n}\|_{C_{p}^{\gamma}[S,T]}+n^{-1})(T-S)^{\varepsilon}.$
@@
$\|X-X^{n}\|_{C_{p}^{\gamma}[S,T]}$ $\leqslant|(X-X^{n})_{S}|+2[X-X^{n}]_{C_{p}^{\gamma}[S,T]}$
$\leqslant|(X-X^{n})_{S}|+C(\|X-X^{n}\|_{C_{p}^{\gamma}[S,T]}+n^{-1})(T-S)^{\varepsilon}.$
@@ 5. Optimality
In this section, we discuss the optimality of the convergence rate $n^{-1}$ obtained in the previous part. By showing that $n(X-X^{n})$ converges to a non-zero limit, we verify that the rate $n^{-1}$ is optimal for the scheme (1.9).
In the following first we present a regularization lemma concerning the solution to (1.3).
Lemma 5.1.
Let $(X_{t})_{t\in[0,1]}$ be the solution to (1.3). Suppose Assumption˜2.1 holds. Then for every $p>1$, for any $f\in{\mathcal{C}}^{1}$, we have
$\Big\|\int_{0}^{\cdot}\nabla f(X_{t})\mathrm{d}t\Big\|_{C_{p}^{1+(\alpha-1)H}[0,1]}\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}},$ (5.1)
where $C$ depends only on $d,p,H,\alpha,\|b\|_{{\mathcal{C}}^{\alpha}}$.
@@
In order to show (5.1), we apply Lemma˜3.1. Let $M=1$, $(s,t)\in[0,1]_{1}^{2}$ and
$A_{s,t}:={\mathbb{E}}^{s-(t-s)}\int_{s}^{t}\nabla f(B_{r}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{r})\mathrm{d}r\text{ where }\varphi:=X-B^{H}.$
We are going to verify (3.1) and (3.2). By (3.7), we see
$A_{s,t}=\int_{s}^{t}\mathcal{P}^{H}_{r-[s-(t-s)]}(\nabla f)({\mathbb{E}}^{s-(t-s)}B_{r}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{r})\mathrm{d}r.$
Then by (3.9), we have
$|A_{s,t}|\leqslant\int_{s}^{t}\|\mathcal{P}^{H}_{r-[s-(t-s)]}(\nabla f)\|_{C^{0}}\mathrm{d}r\leqslant C\int_{s}^{t}\left[r-[s-(t-s)]\right]^{H(\alpha-1)}\|f\|_{C^{\alpha}}\mathrm{d}r.$
Therefore, we have $\|A_{s,t}\|_{L^{p}_{\omega}}\leqslant C\|f\|_{C^{\alpha}}(t-s)^{1-(1-\alpha)H}.$ Then (3.1) holds with $C_{1}=C\|f\|_{C^{\alpha}}$ by the fact that $1-(1-\alpha)H>\frac{1}{2}$.
${\mathbb{E}}^{s_{1}}\delta A_{s,u,t}$ $={\mathbb{E}}^{s_{1}}\bigg[\int_{s_{4}}^{s_{5}}{\mathbb{E}}^{s_{1}}\nabla f(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-{\mathbb{E}}^{s_{3}}\nabla f(B_{r}^{H}+{\mathbb{E}}^{s_{3}}\varphi_{r})\mathrm{d}r\bigg]$
$\quad+{\mathbb{E}}^{s_{1}}\bigg[\int_{s_{5}}^{s_{6}}{\mathbb{E}}^{s_{1}}\nabla f(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-{\mathbb{E}}^{s_{2}}\nabla f(B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r})\mathrm{d}r\bigg]$
$=\int_{s_{4}}^{s_{5}}{\mathbb{E}}^{s_{1}}\bigg[\mathcal{P}^{H}_{r-s_{3}}\nabla f({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-\mathcal{P}^{H}_{r-s_{3}}\nabla f({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{3}}\varphi_{r})\bigg]\mathrm{d}r$
$\quad+\int_{s_{5}}^{s_{6}}{\mathbb{E}}^{s_{1}}\bigg[\mathcal{P}^{H}_{r-s_{2}}\nabla f({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-\mathcal{P}^{H}_{r-s_{2}}\nabla f({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r})\bigg]\mathrm{d}r.$
@@
$\|{\mathbb{E}}^{s_{1}}\delta A_{s,u,t}\|_{L^{p}_{\omega}}$ $\leqslant C\int_{s}^{t}(r-s_{3})^{H(\alpha-2)}\|f\|_{{\mathcal{C}}^{\alpha}}\|{\mathbb{E}}^{s_{1}}[|{\mathbb{E}}^{s_{1}}\varphi_{r}-{\mathbb{E}}^{s_{3}}\varphi_{r}|]\|_{L^{p}_{\omega}}\mathrm{d}r$
$\quad+\int_{u}^{t}(r-s_{2})^{H(\alpha-2)}\|f\|_{{\mathcal{C}}^{\alpha}}\|{\mathbb{E}}^{s_{1}}[|{\mathbb{E}}^{s_{1}}\varphi_{r}-{\mathbb{E}}^{s_{2}}\varphi_{r}|]\|_{L^{p}_{\omega}}\mathrm{d}r$
$\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}|t-s|^{2+2\alpha H-2H}.$
Noticing Assumption˜2.1 implies $2+2\alpha H-2H>1$, we conclude (3.2) holds with $C_{2}=C\|f\|_{{\mathcal{C}}^{\alpha}}$. Moreover, since $f\in C^{1}$, it is obviously to see that the process $\mathcal{A}$ in (3.3) actually is given by
${\mathcal{A}}_{t}=\int_{0}^{t}\nabla f(X_{r})\mathrm{d}r.$
Therefore, by Lemma˜3.1, we have
$\big\|\int_{s}^{t}\nabla f(X_{r})\mathrm{d}r\big\|_{L^{p}}\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}|t-s|^{H(\alpha-1)+1},$
which implies that (5.1) holds. ∎
Due to (5.1), we can already state a result which is analogous to [7, Proposition 2.8].
Proposition 5.2.
Let $(X_{t})_{t\in[0,1]}$ be the solutions to (1.3). Suppose Assumption˜2.1 holds. For every $f\in{\mathcal{C}}^{\alpha}({\mathbb{R}}^{d})$ there exists a $({\mathcal{F}}_{t})_{t\in[0,1]}$-adapted process $({\mathcal{H}}f)^{X}_{t\in[0,1]}\in C_{p}^{1+(\alpha-1)H-}[0,1]$ for every $p\geqslant 2$ such that for $g\in{\mathcal{C}}^{1}$ with probability one,
$({\mathcal{H}}g)^{X}_{t}=\int_{0}^{t}\nabla g(X_{s})\mathrm{d}s,\quad t\in[0,1].$
Proof.
In fact, we notice that for any $f\in{\mathcal{C}}^{\alpha},$ there exists sequence $(f_{n})_{n\in\mathbb{N}}\subset{\mathcal{C}}^{1}$, such that $\lim_{n\to\infty}f_{n}=f$ in ${\mathcal{C}}^{\alpha-}.$ Therefore, by Lemma˜5.1, we have $\big(\int_{0}^{\cdot}\nabla f_{n}(X_{t})\mathrm{d}t\big)_{n\in\mathbb{N}}$ is a Cauchy sequence in $C_{p}^{1+(\alpha-1)H-}[0,1].$ We define ${\mathcal{H}}f=\lim_{n\rightarrow\infty}\int_{0}^{\cdot}\nabla f_{n}(X_{t})\mathrm{d}t$ in $C_{p}^{1+(\alpha-1)H-}[0,1].$
Therefore, for $b\in{\mathcal{C}}^{\alpha}$, we can define $\int_{0}^{\cdot}\nabla b(X_{t})\mathrm{d}t:=\lim_{n\to\infty}\int_{0}^{\cdot}\nabla b_{n}(X_{t})\mathrm{d}t$ in probability. Moreover from Kolmogorov continuity criteria we know that it has a version which has $\theta$-Hölder continuous path for $\theta\in(0,1+(\alpha-1)H)\subset(0,1)$.
Now we are at the position of showing Theorem˜2.4.
Proof of Theorem˜2.4.
Clearly Proposition˜5.2 shows the existence of $({\mathcal{H}}f)^{X}\in C_{p}^{1+(\alpha-1)H-}[0,1]$ which is bounded linear for $f\in{\mathcal{C}}^{\alpha}$, it also implies that the $c$ satisfying (2.7) for given $X$ is well-defined and $\|c\|_{{\mathcal{C}}^{1+(\alpha-1)H-}([0,1])}<\infty,{\mathbb{P}}$-a.s. Indeed from Kolmogorov continuity criteria and the property of Young integral we know that for any possible solution $c$, for any $k_{1}\in(\frac{1}{2},{1+(\alpha-1)H}),k_{2}\in(\frac{1}{2},(1+(\alpha-1)H)\wedge\alpha)=(\frac{1}{2},{1+(\alpha-1)H})$ so that $k_{1}+k_{2}>1$, for any $0\leqslant s\leqslant t\leqslant 1$ and $q>p$, we have
$\|c(t)-c(s)\|_{L^{p}_{\omega}}\leqslant$ $\|\int_{s}^{t}c(r)\mathrm{d}({\mathcal{H}}b)_{r}^{X}\|_{L^{p}_{\omega}}+\frac{1}{2}\|b(X_{t})-b(X_{s})\|_{L^{p}_{\omega}}$
$\leqslant C$ $\big\|\|c\|_{{\mathcal{C}}^{k_{1}}([0,1])}\|({\mathcal{H}}b)^{X}\|_{{\mathcal{C}}^{k_{2}}([0,1])}\big\|_{L^{p}_{\omega}}|t-s|^{k_{2}}+C\|b\|_{{\mathcal{C}}^{\alpha}({\mathbb{R}}^{d})}|t-s|^{\alpha}$
$\leqslant C$ $\|c\|_{L_{\omega}^{q}({\mathcal{C}}^{k_{1}}([0,1]))}\|({\mathcal{H}}b)^{X}\|_{L_{\omega}^{\frac{qp}{q-p}}({\mathcal{C}}^{k_{2}}([0,1]))}|t-s|^{k_{2}}+C\|b\|_{{\mathcal{C}}^{\alpha}({\mathbb{R}}^{d})}|t-s|^{k_{2}}$
$\leqslant C$ $\|b\|_{{\mathcal{C}}^{\alpha}({\mathbb{R}}^{d})}|t-s|^{k_{2}}.$ (5.2)
Then, by Kolmogorov continuity criteria again we get $\mathbb{P}$}-a.s. $\|c\|_{{\mathcal{C}}^{1+(\alpha-1)H-}}<\infty$. Notice that (2.7) is a linear ODE in the sense of Young integral for given $X$, standard fixed point arguments gives us the existence and uniqueness of $c\in{\mathcal{C}}^{1+(\alpha-1)H-}([0,1])$ satisfying (2.7).
Therefore we only need to show that (2.6) holds for such $c$ and for $b\in{\mathcal{C}}^{\alpha}$, $\alpha<1$.
Denote $c^{n}:=n(X-X^{n})$. Theorem˜2.2 implies that $\sup_{n}\|c_{n}\|_{{\mathcal{C}}_{p}^{1+(\alpha-1)H-}[0,1]}<\infty,$ that is to say, for $\gamma<1+(\alpha-1)H$, from Kolmogorov continuity criteria, $\sup_{n}\|c_{n}\|_{L^{p}(\Omega,{\mathcal{C}}^{\gamma})}<\infty$, hence for ${\mathcal{L}}^{n}:=(X,X^{n},c^{n},c,c^{n}-c,B^{H})$, $\sup_{n}\|{\mathcal{L}}^{n}\|_{L^{p}(\Omega,({\mathcal{C}}^{\gamma})^{6})}<\infty.$
By Sobolve embedding we know that there exists large $q$ and $\gamma_{1}<\gamma_{2}<\gamma$ so that compactly
${\mathcal{C}}^{\gamma}([0,1],{\mathbb{R}}^{d})\subset\subset W_{q}^{\gamma_{2}}([0,1],{\mathbb{R}}^{d})\subset\subset{\mathcal{C}}^{\gamma_{1}}([0,1],{\mathbb{R}}^{d}),$ (5.3)
hence the laws of $({\mathcal{L}}^{n})_{n\in{\mathbb{N}}}$ are tight on $W_{q}^{\gamma_{2}}([0,1],{\mathbb{R}}^{d})^{6}$ which is Polish and
$\sup_{n}\||{\mathcal{L}}^{n}|\|_{L^{p}(\Omega,W_{q}^{\gamma_{2}}([0,1],{\mathbb{R}}^{d})^{6})}<\infty.$ (5.4)
[…]
and
$\hat{\mathbb{P}}-a.s.,\quad(\hat{\mathcal{L}}^{n})_{n\in{\mathbb{N}}_{1}}\quad\text{converges to }\quad{\mathcal{L}}\quad\text{ in }W_{q}^{\gamma_{2}}([0,1],{\mathbb{R}}^{d})^{6}.$ (5.6)
Let $(\hat{\mathcal{F}}_{t})_{t\geqslant 0}$ be the augmentation of the filtration generated by $(\hat{X},\hat{X}^{n},\hat{B}^{H,n})$. (5.5), the continuity of $b$ and the strong well-posedness of (1.3) imply
$\tilde{X}_{t}$ $=x_{0}+\int_{0}^{t}b(\tilde{X}_{s})\mathrm{d}s+\hat{B}_{t}^{H},\quad\tilde{X}_{t}^{k}=x_{0}+\int_{0}^{t}b(\tilde{X}_{s}^{k})\mathrm{d}s+\hat{B}_{t}^{H,k},$
$\hat{X}_{t}^{k,n}$ $=x_{0}+\int_{0}^{t}b(\hat{X}_{k_{n}(s)}^{k,n})\mathrm{d}s+\hat{B}_{t}^{H,k},\quad\hat{c}^{n}=n(\tilde{X}^{n}-\hat{X}^{n,n}),$
$\tilde{c}^{n}(t)$ $=\int_{0}^{t}\tilde{c}^{n}(s)\mathrm{d}({\mathcal{H}}b)^{\tilde{X}_{s}^{n}}+\frac{1}{2}(b(\tilde{X}_{t}^{n})-b(x_{0})),\quad t\in[0,1].$
Following from (5.5) which implies the uniform integrability, (5.6) and Vitali’s theorem, for all $p\geqslant 1$ we have
$\lim_{n\rightarrow\infty}\big\|\hat{\mathcal{L}}^{n}-{\mathcal{L}}\big\|_{L^{p}(\hat{\Omega},W_{q}^{\gamma_{2}}([0,1],{\mathbb{R}}^{d})^{6})}=0.$ (5.7)
At this point we can also claim that $\tilde{X}=\hat{X} ^\hat{\mathbb{P}}$-a.s. from Theorem˜2.2 and the strong well-posedness of (1.3).
Notice that for any $b\in{\mathcal{C}}^{\alpha}$, there exists a sequence $\{b_{n}\}_{n\in\mathbb{N}}\subset{\mathcal{C}}^{1}$ such that $\lim_{n\to\infty}b_{n}=b$ in ${\mathcal{C}}^{\alpha-}$, then for $m\in{\mathbb{N}}$
$\hat{c}^{n}(t)=$ $n(\tilde{X}_{t}^{n}-\hat{X}_{t}^{n,n})=n\left(\int_{0}^{t}b(\tilde{X}_{s}^{n})\mathrm{d}s-\int_{0}^{t}b(\hat{X}_{k_{n}(s)}^{n,n})\mathrm{d}s\right)$
$=$ $n\int_{0}^{t}(b-b_{m})(\tilde{X}_{s}^{n})-(b-b_{m})(\hat{X}_{k_{n}(s)}^{n,n})\mathrm{d}s+n\int_{0}^{t}b_{m}(\tilde{X}_{r}^{n})-b_{m}(\hat{X}_{r}^{n,n})\mathrm{d}s$
$+n\int_{0}^{t}b_{m}(\hat{X}_{r}^{n,n})-b_{m}(\hat{X}_{k_{n}(s)}^{n,n})\mathrm{d}s=:J_{t}^{m,n}+K_{t}^{m,n}+L_{t}^{m,n}.$ (5.8)
Following from Lemma˜4.2, Lemma˜4.3 and Lemma˜4.4 we get
$\big\|\sup_{t\in[0,1]}|J_{t}^{m,n}|\big\|_{L_{\omega}^{p}}\lesssim\|b-b_{m}\|_{{\mathcal{C}}^{\alpha-}};$ (5.9)
for $K^{m,n}$
$K_{t}^{m,n}=\int_{0}^{t}\int_{0}^{1}\nabla b_{m}(\theta\tilde{X}_{r}^{n}+(1-\theta)\hat{X}_{r}^{n,n})\cdot\hat{c}_{r}^{n}\mathrm{d}\theta\mathrm{d}r,$
since $^\hat{\mathbb{P}}$-a.s. $\hat{c}^{n}$ converges to $\hat{c}$, $\hat{X}^{n,n}$ converges to $\hat{X}$, and $\tilde{X}^{n}$ converges to $\tilde{X}=\hat{X}$, $^\hat{\mathbb{P}}$-a.s. we have
$\lim_{n\rightarrow\infty}K_{t}^{m,n}=\int_{0}^{t}\nabla b_{m}(\hat{X}_{r})\cdot\hat{c}_{r}\mathrm{d}r=\int_{0}^{t}\hat{c}_{r}\mathrm{d}({\mathcal{H}}b_{m})_{r}^{\hat{X}},\quad\hat{\mathbb{P}}-a.s.$ (5.10)
@@
Concerning $L^{m,n}$, following from (1.9) and the fact a.s. $B^{H}\in{\mathcal{C}}^{H-\varepsilon}([0,1])$ for sufficiently small $\varepsilon>0$ by Kolmogorov continuity criteria, which means there exists a small enough $\varepsilon^{\prime}>0$ so that a.s. $B_{H}\in{\mathcal{C}}^{1+\varepsilon^{\prime}}$, we have a.s. (denote $\{c\}:=c-\lfloor c\rfloor$ for $c\in{\mathbb{R}}_{+}$)
$n(\hat{X}_{r}^{n}-\hat{X}_{k_{n}(r)}^{n,n})$ $=n(r-k_{n}(r))b(\hat{X}_{k_{n}(r)}^{n,n})+n(\hat{B}_{r}^{H,n}-\hat{B}_{k_{n}(r)}^{H,n})$
$=\{nr\}b(\hat{X}_{k_{n}(r)}^{n,n})+\{nr\}(\hat{B}_{r}^{H,n})^{\prime}+o(n^{-\varepsilon^{\prime}});$